OpenStudy (anonymous):

determine the type and number of solutions of x^2-8x+16=0

OpenStudy (anonymous):

do you need it to be solved?

OpenStudy (anonymous):

x2−8x+16=0 Step 1: Factor left side of equation. (x−4)(x−4)=0 Step 2: Set factors equal to 0. x−4=0 or x−4=0 x=4 Answer: x=4

OpenStudy (dayakar):

equal roots